the four angles of a quadrilateral have the same constant difference between hem. if the smallest angle is 75 deg, find the measure of the second largest angle
Answer by Edwin McCravy(20060)

Let the constant difference between the angles be c

smallest angle = 75°
next to smallest angle = 75°+c
next to largest angle = 75°+c+c = 75°+2c
largest angle = 75°+2c+c = 75°+3c

The sum of the four angles of any quadrilateral is 360°

75° + (75°+c) + (75°+2c) + (75°+3c) = 360°
      75° + 75°+c + 75°+2c + 75°+3c = 360°
                          300° + 6c = 360°
                                 6c = 60°
                                  c = 30°

smallest angle = 75° 
next to smallest angle = 75°+c = 75°+30° = 105°
next to largest angle = 75°+2c = 75°+2(30°) = 75°+60° = 135° 
largest angle = 75°+2c+c = 75°+3c = 75°+3(30°) = 75°+90° = 165°

[Note:  You may point out to your teacher that this is a quadrilateral,
not a trapezium, because a trapezium has two parallel sides, and
therefore it must have two pairs of supplementary angles, and there 
is only one pair of supplementary angles, 75² and 105².  That's why
I changed the word "trapezium" to "quadrilateral".]

Edwin